A proportional relationship between the number of pounds of potatoes (x) and the price in dollars (y) is graphed, and the ordered pair (4,3) is on the graphed line. What is the price of 1 pound of potatoes? Show your work, including the proportion you used to determine the price. What does the ordered pair (8,6) on the graph represent? Explain in words.
Answer by TimothyLamb(4379)

dollars/pounds = 3/4
pounds = 1
dollars/1 = 3/4
dollars = 1*3/4
dollars = 3/4
dollars = 0.75
---
(8,6)
8 pounds of potatoes costs 6 dollars
---
Free algebra tutoring live chat:
https://sooeet.com/chat.php?gn=algebra
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations with quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php
---